# Nonlinear Permuted Granger Causality

### v1

@@ 1.3 Overview
Rather than comparison of the inherently unequal restricted and unrestricted model errors, we focus out-of-sample predictability by implementing a permutation structure. There is precedence for the use of permutation-type procedures on general linear models, and their asymptotics are well studied in literature (Anderson and Legendre, 1999; Anderson and Robinson, 2001; DiCiccio and Romano, 2017; Winkler et al., 2014). Nauta et al. (2019) employ this type of importance measure after their complicated CNN screening procedure, and we argue for its widespread use as a decision framework in Granger causal inference. Importance of a chronologically ordered variable can be interpreted as “causal” (predictive) effect, and the strategy builds on the concept of exchangeability (like the directed graph method of Caron and Fox (2017)), where if $\tilde{\mathbf{X}}$ is a random permutation of the rows of $\mathbf{X}$, $\mathbf{Y}_{<t+1}\perp\mathbf{X}$ given $\mathbf{Y}_{<t}$ implies $\mathbf{Y}_{<t+1}\perp\tilde{\mathbf{X}}$ given $\mathbf{Y}_{<t}$, but the converse is not always true (Van der Laan, 2006).
To investigate if $\mathbf{X}$ Granger causes $\mathbf{Y}$, we propose the following definition pair, adjusting Definitions 2 and 3 to a permutation structure for the covariate matrix. The unrestricted and restricted models are replaced by a null model and a permuted model.
Definition 4a.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is not conditionally Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the relevant history of the response $\mathbf{Y}_{<t}$ if and only if $\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\overset{d}{=}\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z},\tilde{\mathbf{X}}$, where $\tilde{\mathbf{X}}$ is a copy of $\mathbf{X}$ with the time axis (rows) randomly permuted.
Definition 4b.
@@
The permutations of $\mathbf{X}$ break the potential dependence structure with the response while retaining the intradependence of the covariates, as information in a given row across columns remains intact. The restructuring of the Granger causal framework allows for use of out-of-sample estimated prediction errors, corrects the imbalance of comparison between restricted and unrestricted, and presents a clear path to account for multiple testing, aligning the methodology closer to its inferential utility.
In Section 2, we outline the nonlinear permuted Granger causal methodology (NPGC) as an alternative framework to penalized optimization approaches, and we discuss the theoretical results in Section 3. Section 4 presents performance for the permutation method for simulated data, Section 5 applies NPGC to publicly available whole-cell recordings from the primary auditory cortex of an anesthetized rat in response to natural sound fragments, and Section 6 concludes the manuscript.
@@ 2 Methodology
Suppose observed realizations of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ arise from the set of all potential realizations $\omega\in\Omega$. We define $\Omega_{\text{obs}}$ as the size $\varphi$ set of observations, $\Omega_{\text{obs}}=\left\{1,\ldots,\varphi\right\}\subseteq\Omega$, and note that usually $\varphi=1$. Instances for $\varphi>1$ may occur with repeated trials of a controlled experiment. For simplicity of the original presentation, we omit the subscript $\omega$ notation specifying an observed realization until the end of this subsection.
With familiarity of a dataset, a researcher selects $\gamma$ lagged values of $\mathbf{Y}$ that serve as a representative history, $\mathbf{Y}_{\text{lag}}\in\mathbb{R}^{(T+\gamma)\times\gamma d}$ (corresponding to $\mathbf{Y}_{<t}$ in Definitions 4a and 4b), and assume that an appropriate number of lags is selected such that autocorrelation in $\mathbf{Y}$ is fully explained across all potential realizations. Selection of the truncation lag $\gamma$ is outside the scope of this work, and we refer the reader to Ng and Perron (2001), Ivanov and Kilian (2005), Shojaie and Michailidis (2010), and Nicholson et al. (2017) for detailed discussions. With a finite data length $T+\gamma$, this process restricts the usable portion of $\mathbf{X}$, $\mathbf{Y}$, and any additional covariates $\mathbf{Z}$ to the last $T$ rows. All columns are standardized for consistent behavior in the featurization process with an activation function (Goodfellow et al., 2016).
@@
After permutation, we compile the predictor matrices $\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\in\mathbb{R}^{T\times(1+\gamma d+q+p)}$, and set the nonlinear activation function $g=\tanh$ in the functional transformation $\Psi$ of Equation 1. With the inclusion of the intercept term, $\Psi$ becomes
$\mathbf{H}_{m}$ $=\Psi\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\right)=\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\mathbf{W}\right),$
$\text{where}\quad\mathbf{W}$ $=(w_{ij})\in\mathbb{R}^{(1+\gamma d+q+p)\times N},\quad w_{ij}\sim\mathcal{N}(0,1).$ (4)
The transformation matrix is generated via random initialization as indicated. We account for added uncertainty in these random models by generating several initializations $\mathbf{W}_{r}$ for $r=1,\ldots,\mathscr{R}$ and extracting the aggregate behavior. We can write the models in terms of the original ($m=1$) and permuted ($m=2,\ldots,M$) feature spaces: $\mathbf{Y}=\mathbf{H}_{m,r}\mathbf{B}_{m,r}+\mathbf{U}_{m,r}$ (where $\mathbf{B}_{m,r}$ corresponds to $\mathbf{W}^{L}$ in Equation 3). The exact formulation of $\Psi$ in Equation 4 is not the main focus of this manuscript and the dimension of the feature space $N$ is fixed for direct comparison to other methods. For demonstration of our method, we select a simple, widely-applicable model agnostic structure, and discuss potential adaptations in the supplement, including a cross validation approach to select an optimal dimension.
Define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the prediction for $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given the relevant history of the response $\mathbf{Y}_{\text{lag}}$, the additional variables $\mathbf{Z}$, and the permuted covariate set $\tilde{\mathbf{X}}_{m}$. We denote $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$ as the corresponding parameter over all potential realizations $\omega\in\Omega$. Variation in the estimate arises from potential realizations of the data $\omega\in\Omega$ and via random initializations approximating the nonlinear functional form. For a given data realization $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, and under the true functional form, we define the specific covariance matrix of the prediction $\bm{\Sigma}_{m,\omega}$; for random initialization $r$, we define the covariance matrix of prediction $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$. We estimate the out-of-sample variation parameter $\vartheta_{m}$ for each permutation with a cross validation approach.
@@ 2.1 Estimating Granger Causal Influence
A sufficiently large value featurization dimension $N$ is chosen to “linearize” any existing functional relationship, but not so large that the network is able to memorize inputs or fabricate dependence between the permuted data and a response. We implicitly assume the existence of some nonlinear functional relationship between a covariate set and a response will be “easier” for an artificial network to learn than random matching of inputs to outputs in the permuted data, and propose an exact form of this condition in Section 3. The data is split into $K$ sets for model computation and testing, and we define the number of observations in each set $k=1,\ldots,K$ as $T_{k}=\lfloor T/K\rfloor+\mathbf{1}\left\{\left(T\bmod K\right)\geq k\right\}$.
We generate $r=1,\ldots,\mathscr{R}$ featurizations under the form of Equation 4. The model matrices $\mathbf{W}_{r}$ are held fixed over all permutations $m$ and observations $\omega\in\Omega_{\text{obs}}$ for a consistent error estimation framework. For each permutation, with $m=1$ corresponding to the original data, the predictor matrices are projected into the respective feature spaces, and we can write the residuals for test set $k$ as
$\mathbf{R}_{\omega,m,r,k}$ $=\mathbf{H}_{\omega,m,r,k}\left(\mathbf{H}_{\omega,m,r,-k}^{\prime}\mathbf{H}_{\omega,m,r,-k}\right)^{-1}\mathbf{H}_{\omega,m,r,-k}^{\prime}\mathbf{Y}_{\omega,-k},-\mathbf{Y}_{\omega,k}$ (5)
where the training data (subscript k-k) excludes set $k$.
These out-of-sample prediction residuals from the test set align closer to the original definition of predictive ability in Granger causal inference than the in-sample model variation. This distinction is especially important with the use of artificial networks and the ability to learn any arbitrary, data-specific dependence structure. We obtain the estimate for the out-of-sample variation in prediction residuals $\hat{\vartheta}_{m}$ with Equation 6.
$\hat{\vartheta}_{m}$ $=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}_{\omega,m,r,k}^{\prime}\mathbf{R}_{\omega,m,r,k}\right)$ (6)
Using the permutation structure, we approximate a null distribution of variation from each model $m=1,\ldots,M$. These random permutations of the covariate set $\mathbf{X}$ break the potential dependence structure present in the form of some predictive relationship with the response $\mathbf{Y}$. Under the null hypothesis, we can view the original data as one of $M$ random permutations. We expect that the original “permutation” $\tilde{\mathbf{X}}_{1}=\mathbf{X}$ will exhibit similar properties to $\tilde{\mathbf{X}}_{m}$ for $m=2,\ldots,M$, or if the time observations $\mathbf{X}$ are exchangeable for prediction of $\mathbf{Y}$, the conditional distribution of the prediction does not change. The variation estimates are drawn from a distribution of all possible permutations,
@@
and we expect $\hat{\vartheta}_{1}$ to fall above some lower tail portion. This leads to an approximate null distribution where the estimate $\hat{\vartheta}_{1}$ is at quantile $\hat{Q}_{M}$ of the empirical distribution $\hat{\mathcal{H}}_{M}(s)$, defined in Equation 8, formed from a subsample of size $M\leq T!$. A decision rule is formulated from comparison to a chosen level of test $\alpha$.
@@
Rejection of the null hypothesis in this framework, $\hat{Q}_{M}\leq\alpha$, presents evidence for $\mathbf{X}$ as Granger causal of $\mathbf{Y}$ conditional on the additional variables $\mathbf{Z}$ and the relevant history of the response $\mathbf{Y}_{\text{lag}}$. For the case when $\varphi=1$, we also clarify that this conclusion is conditional on error free observation of the dataset. Inferential results must either include this assumption, or the scope narrowed to the observation $\omega$.
The full process is given in the form of pseudocode in the supplementary material. We explicitly outline the theoretical behavior of these estimates and the development of the underlying framework in Section 3.
@@ 3 Theory
We define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the predictive ability of permutation $m$, and the quantity $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$. For each potential realization of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, $\omega\in\Omega$, we define the realization-specific covariance matrix $\bm{\Sigma}_{m,\omega}$ that is drawn from some distribution with expectation $\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]=\vartheta_{m}$ and variance $\tau_{\omega}^{2}<\infty$ that is constant over all permutations. Each random initialization for the featurization process $r=1,\ldots,\mathscr{R}$ produces $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$, where $\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and $\text{Var}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\tau_{r}^{2}<\infty$.
Using the permutation structure, we evaluate the null and permuted models with the out-of-sample prediction residuals, shown in Equation 5, and define the estimate for total variation in Equation 6. Under the null hypothesis when the conditional distribution of the prediction is invariant to permutation of the covariate set $\mathbf{X}$, $\vartheta_{1}=\vartheta_{2}=\cdots=\vartheta_{T!}=\vartheta$, and we write the null and alternative hypotheses below.
@@
We test the null hypothesis in Equation 10 using the sample quantile $\hat{Q}_{M}$ from the empirical distribution $\hat{\mathcal{H}}_{M}(s)$ defined in Equations 8 and 9, and outline theoretical results under a set of mild conditions that are comparable to those found in relevant literature. Four additional conditions in the supplement provide regularity to the featurization process.
Condition 1.
The data is continuous and stationary, and the discrete, regular sampling grid $t=1,\ldots,T$ is sufficiently fine to capture any potential functional dependence in the variable matrices $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for any realization $\omega\in\Omega$.
Condition 2.
For all realizations $\omega\in\Omega$, and random initializations $\mathbf{W}_{r}\in\mathcal{W}$, $r=1,\ldots,\mathscr{R}$, the quantities $\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}$ are drawn i.i.d. from continuous distributions with $\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and constant variance across permutations and potential realizations $\text{Var}\left(\text{tr}\left[\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)=\tau_{r}^{2}<\infty$. Similarly, $\bm{\Sigma}_{m,\omega}$ are i.i.d. from continuous distributions with expectation $\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]=\vartheta_{m}$ and constant variance across permutations $\text{Var}\left(\text{tr}\left[\bm{\Sigma}_{m,\omega}\right]\right)=\tau_{\omega}^{2}<\infty$.
Condition 3.
The relevant history of the response, $\mathbf{Y}_{<t,\omega}=\mathbf{Y}_{\text{lag},\omega}\in\mathbb{R}^{T\times\gamma d}$ is appropriately chosen such that the model errors $\mathbf{U}_{m,\omega,r}$ are independent, or $\mathbf{u}_{m,\omega,r,t}\perp\mathbf{u}_{m,\omega,r,t^{\prime}}$ for any realization $\omega\in\Omega$, permutation $m=1,\ldots,M$, initialization $r=1,\ldots,\mathscr{R}$, and time point $t=1,\ldots,T$ where $t^{\prime}\neq t$. Further, the model errors follow multivariate normal distributions with mean zero and constant variation, leading to the result

### v2

@@ 1.2 Nonlinear Adaptations
Rather than comparison of the inherently unequal restricted and unrestricted model errors, focus is shifted to out-of-sample predictability by implementing a permutation structure. There is precedence for the use of permutation-type procedures on general linear models, and their asymptotics are well studied in literature (Anderson and Legendre, 1999; Anderson and Robinson, 2001; DiCiccio and Romano, 2017; Winkler et al., 2014). Nauta et al. (2019) employ this type of importance measure after their complicated CNN screening procedure, this work explicitly defines the methodology for its widespread use as a decision framework in Granger causal inference. Importance of a chronologically ordered variable can be interpreted as “causal” (predictive) effect, and the strategy builds on the concept of exchangeability (like the directed graph method of Caron and Fox (2017)), where if $\tilde{\mathbf{X}}$ is a random permutation of the rows of $\mathbf{X}$, $\mathbf{Y}_{<t+1}\perp\mathbf{X}$ given $\mathbf{Y}_{<t}$ implies $\mathbf{Y}_{<t+1}\perp\tilde{\mathbf{X}}$ given $\mathbf{Y}_{<t}$, but the converse is not always true (Van der Laan, 2006).
@@
$\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\mathbf{X}_{<t}\overset{d}{=}\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\tilde{\mathbf{X}}_{<t}.$ (8)
{subdefinition}
Equivalent to Definition 1.2, $\mathbf{x}_{t}$ is conditionally Granger causal for $\mathbf{y}_{t}$ given $\mathbf{z}_{t}$ and the relevant history of the response $\mathbf{Y}_{<t}$ if
$\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\mathbf{X}_{<t}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\tilde{\mathbf{X}}_{<t}\right)\right].$ (9)
The permutations of $\mathbf{X}_{<t}$ break the potential dependence structure with the response while retaining the intradependence of the covariates, as information in a given row across columns remains intact. The restructuring of the Granger causal framework allows for use of out-of-sample estimated prediction errors, corrects the imbalance of comparison between restricted and unrestricted, and presents a clear path to account for multiple testing, aligning the methodology closer to its inferential utility.
@@ 2 Methodology
Suppose observed realizations of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ arise from the set of all potential realizations $\omega\in\Omega$. Define $\Omega_{\text{obs}}$ as the size $\varphi$ set of observations, $\Omega_{\text{obs}}=\left\{1,\ldots,\varphi\right\}\subseteq\Omega$, and note that usually $\varphi=1$. Instances for $\varphi>1$ may occur with repeated trials of a controlled experiment. For simplicity of the original presentation, the subscript $\omega$ notation specifying an observed realization is omitted until the end of this subsection.
2.1 Structure
Nonlinear functional dependence in the data is captured with feed forward networks (FNNs) like Equations 5 and 7, where the depth $L=1$, and $\mathbf{W}^{0}$ and $\mathbf{b}^{0}$ are randomly generated. The exact formulation of this transformation to a representative space (akin to $\Psi$ in Section 1.2) is not the main focus of this work, and the dimension of the feature space $N$ is fixed for direct comparison to other methods. For demonstration of NPGC, a simple model agnostic structure was selected. Other, more targeted featurization strategies will likely more effectively describe data specific dependence.
With familiarity of a dataset, a researcher selects $\gamma$ lagged values of $\mathbf{y}_{t}$ that serve as a representative history, $\mathbf{Y}_{\text{lag}}\in\mathbb{R}^{(T+\gamma)\times\gamma d}$ (corresponding to $\mathbf{Y}_{<t}$ in Definitions 1.2 and 1.2), and assume that an appropriate number of lags is selected such that autocorrelation in $\mathbf{y}_{t}$ is fully explained across all potential realizations. Selection of the truncation lag $\gamma$ is outside the scope of this work; Ng and Perron (2001); Ivanov and Kilian (2005); Shojaie and Michailidis (2010); Nicholson et al. (2017) provide detailed discussions. With a finite data length $T+\gamma$, this process restricts the usable portion of $\mathbf{X}$, $\mathbf{Y}$, and any additional covariates $\mathbf{Z}$ to the last $T$ rows. All columns (individual variables) are standardized for consistent behavior in a random FNN featurization process with an activation function (Goodfellow et al., 2016).
@@ 2.1 Structure
After permutation, the predictor matrices $\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\in\mathbb{R}^{T\times(1+\gamma d+q+p)}$ are compiled, and the FNN is rewritten to the structure in Equation 10 (with activation function $g=\tanh$). $\mathbf{W}^{0}$ and $\mathbf{b}^{0}$ are combined into a single parameter matrix $\mathbf{W}\in\mathbb{R}^{(1+\gamma d+q+p)\times N}$ after inclusion of the intercept term in the predictor matrices, and each matrix entry is an independent Gaussian realization $w_{ij}\sim\mathcal{N}(0,1)$.
$\mathbf{H}_{m}$ $=g\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\mathbf{W}\right)=\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\mathbf{W}\right)$ (10)
Added uncertainty arising from the random generation is mitigated through several featurizations, $\mathbf{W}_{r}$ for $r=1,\ldots,\mathscr{R}$, and extracting the aggregate behavior. The models can be written in terms of the original ($m=1$) and permuted ($m=2,\ldots,M$) feature spaces, where $\mathbf{U}_{m,r}$ is the variation in $\mathbf{Y}$ not captured by the functional relationship with the feature space $\mathbf{H}_{m,r}$.
$\mathbf{Y}=\mathbf{H}_{m,r}\mathbf{W}^{L}_{m,r}+\mathbf{U}_{m,r}$ (11)
Define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the prediction for $\mathbf{Y}$ given the relevant history of the response $\mathbf{Y}_{\text{lag}}$, the additional variables $\mathbf{Z}$, and the permuted covariate set $\tilde{\mathbf{X}}_{m}$. Denote $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$ as the corresponding parameter over all potential realizations $\omega\in\Omega$. Variation in the estimate arises from potential realizations of the data $\omega\in\Omega$ and via randomly generated FNNs approximating the nonlinear functional form. For a given data realization $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, and under the true functional form $f$, define the specific covariance matrix of the prediction $\bm{\Sigma}_{m,\omega}$. For random featurization $r$, define the covariance matrix of prediction $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$. The out-of-sample variation parameter $\vartheta_{m}$ is estimated for each permutation via a cross-validation approach.
@@ 2.2 Estimating Granger Causal Influence
A sufficiently large value featurization dimension $N$ is chosen to “linearize” any existing functional relationship, but not so large that the network is able to memorize inputs or fabricate dependence between the permuted data and a response. This implicitly assumes the existence of some nonlinear functional relationship between a covariate set and a response will be “easier” for an ANN to learn than random matching of inputs to outputs in the permuted data, and an exact form of this condition is proposed in Section 3. The data is split into $K$ sets for model computation and testing, and define the number of observations in each set $k=1,\ldots,K$ as $T_{k}=\lfloor T/K\rfloor+\mathbf{1}\left\{\left(T\bmod K\right)\geq k\right\}$.
Under the form of Equation 10, several random FNNs are generated $r=1,\ldots,\mathscr{R}$. The model matrices $\mathbf{W}_{r}$ are held fixed over all permutations $m$ and observations $\omega\in\Omega_{\text{obs}}$ for a consistent error estimation framework. For each permutation, with $m=1$ corresponding to the original data, the predictor matrices are projected into the respective feature spaces, and the residuals for test set $k$ can be written as in Equation 12, where the where the training data (subscript k-k) excludes set $k$.
$\mathbf{R}_{m,\omega,r,k}$ $=\mathbf{H}_{m,\omega,r,k}\left(\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right)^{-1}\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{Y}_{\omega,-k},-\mathbf{Y}_{\omega,k}$ (12)
The out-of-sample prediction residuals $\mathbf{R}$ from the test set align closer to the original definition of predictive ability in Granger causal inference than the in-sample model variation. This distinction is especially important with the use of ANNs and the ability to learn any arbitrary, data-specific dependence structure. The estimate for the out-of-sample variation in prediction residuals $\hat{\vartheta}_{m}$ is shown in Equation 13.
$\hat{\vartheta}_{m}$ $=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)$ (13)
A null distribution of variation from each model $m=1,\ldots,M$ is approximated from the permutation structure. The random permutations of the covariate set $\mathbf{X}$ break the potential dependence structure present in the form of some predictive relationship with the response $\mathbf{Y}$. Under the null hypothesis, the original data is viewed as one of $M$ random permutations, and the original “permutation” $\tilde{\mathbf{X}}_{1}=\mathbf{X}$ should exhibit similar properties to $\tilde{\mathbf{X}}_{m}$ for $m=2,\ldots,M$. Analogously, if the time observations $\mathbf{X}$ are exchangeable for prediction of $\mathbf{Y}$, the conditional distribution of the prediction will not change.
The variation estimates are drawn from the distribution of all possible permutations in Equation 14, and $\hat{\vartheta}_{1}$ is expected to fall above some lower tail portion.
@@
This leads to an approximate null distribution where the estimate $\hat{\vartheta}_{1}$ is at quantile $\hat{Q}_{M}$ of the empirical distribution $\hat{\mathcal{H}}_{M}(s)$, defined in Equation 15, formed from a subsample of size $M\leq T!$. A decision rule is formulated from comparison to a chosen level of test $\alpha$.
@@
Rejection of the null hypothesis in this framework, $\hat{Q}_{M}\leq\alpha$, presents evidence for $\mathbf{X}$ as Granger causal of $\mathbf{Y}$ conditional on the additional variables $\mathbf{Z}$ and the relevant history of the response $\mathbf{Y}_{\text{lag}}$. For the case when $\varphi=1$, this conclusion is conditional on error free observation of the dataset. Inferential results must either include this assumption, or the scope narrowed to the specific observation $\omega$. The full algorithmic process is shown in the supplement, and an explicit outline of the theoretical behavior of these estimates and development of the underlying framework is given in Section 3.
@@ 3 Theory
Define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the predictive ability of permutation $m$, and the quantity $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$. For each potential realization of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, $\omega\in\Omega$, define the realization-specific covariance matrix $\bm{\Sigma}_{m,\omega}$ drawn from some distribution with expectation $\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]=\vartheta_{m}$ and variance $\tau_{\omega}^{2}<\infty$ that is constant over all permutations. Each random generated FNN for the featurization process $r=1,\ldots,\mathscr{R}$ produces $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$, where $\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and $\text{Var}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\tau_{r}^{2}<\infty$.
The null and permuted models are evaluated with the out-of-sample prediction residuals, shown in Equation 12, and define the estimate for total variation as in Equation 13. Under the null hypothesis when the conditional distribution of the prediction is invariant to permutation of the covariate set $\mathbf{X}$, $\vartheta_{1}=\vartheta_{2}=\cdots=\vartheta_{T!}=\vartheta$, leading to the null and alternative hypotheses in Equation 17.
@@
The null hypothesis is tested using the sample quantile $\hat{Q}_{M}$ from the empirical distribution $\hat{\mathcal{H}}_{M}(s)$ defined in Equations 15 and 16. Theoretical results in this section rely on a set of three mild conditions comparable to those found in relevant literature. Four additional conditions provide regularity to the featurization process. {condition} The data is continuous and stationary, and the discrete, regular sampling grid $t=1,\ldots,T$ is sufficiently fine to capture any potential functional dependence in the variable matrices $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for any realization $\omega\in\Omega$. {condition} For all realizations $\omega\in\Omega$, and random generated FNNs $\mathbf{W}_{r}\in\mathcal{W}$, $r=1,\ldots,\mathscr{R}$, the quantities $\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}$ are independently drawn from continuous distributions with defined expectation in Equation 18 and constant, finite variance across all permutations and potential realizations of the data.
$\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]$ $=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ (18)
$\text{Var}\left(\text{tr}\left[\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)$ $=\tau_{r}^{2}<\infty$ (19)
Similarly, $\bm{\Sigma}_{m,\omega}$ are independently drawn from continuous distributions with expectation in Equation 20 and constant, finite variance across permutations.
$\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]$ $=\vartheta_{m}$ (20)
$\text{Var}\left(\text{tr}\left[\bm{\Sigma}_{m,\omega}\right]\right)$ $=\tau_{\omega}^{2}<\infty$ (21)
{condition}
The relevant history of the response, $\mathbf{Y}_{<t,\omega}=\mathbf{Y}_{\text{lag},\omega}\in\mathbb{R}^{T\times\gamma d}$ is appropriately chosen such that the model errors $\mathbf{U}_{m,\omega,r}$ are independent, or $\mathbf{u}_{m,\omega,r,t}\perp\mathbf{u}_{m,\omega,r,t^{\prime}}$ for any realization $\omega\in\Omega$, permutation $m=1,\ldots,M$, featurization $r=1,\ldots,\mathscr{R}$, and time point $t=1,\ldots,T$ where $t^{\prime}\neq t$. Further, the model errors follow multivariate normal distributions with mean zero and constant variation, leading to the result